# Dalitz Plot Kinematics for a Lorentz-Violating Three-Body Decay

### v1

@@ 1 Introduction
The standard model of physics describes a theory that is invariant under rotations and Lorentz transformations, but theories with Lorentz symmetry breaking frames are also of interest, since quantum theories of gravitation may not have the same symmetries as the low-energy effective theories that we see in everyday operation. Theories of Planck-scale physics like strings, loop quantum gravity, noncommutative spacetime structures, spacetime foam, nontrivial spacetime topologies, and others already predict or strongly suggest Lorentz violation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. The same tools used to study the standard model may also be used to analyze more general theories in which Lorentz symmetry is weakly broken. The general local field theory that describes Lorentz-violating modifications to processes involving known standard model fields has been laid out in detail [15, 16]. This theory, which is called the standard model extension (SME), is also the general theory for describing local CPT violation involving those known fields (although, while the SME may be extended to include gravity [17], gravitational CPT violation is a more complicated matter [18, 19, 20]); the reason is that a local, stable, unitary, CPT-violating quantum field theory is also necessarily Lorentz violating [21]. The SME was constructed precisely to provide a systematic approach, via an effective field theory framework for understanding searches for violations of the fundamental Lorentz and CPT symmetries.
@@
As the SME is an effective field theory, it contains towers of operators of increasing mass dimension, with the contributions from higher-dimensional operators assumed to be further suppressed under most circumstances. The minimal SME is a restricted subtheory of the SME—that which is expected to be renormalizable, because it satisfies the usual conditions that ensure renormalizability for the standard model, by containing only the finite number of local, Hermitian, gauge-invariant operators which can be built out the known standard model fermion and boson field that are of dimension four or less. The terms in the minimal SME Lagrange density are, in structure, very similar to those appearing the usual standard model; the key difference that the additional Lorentz-violating operators each have one or more uncontracted Lorentz indices. The coefficients multiplying those operators then become observable as components of preferred background vectors or tensors. In many situations, the minimal SME is the most natural test theory for analyzing the results of experimental searches for CPT or Lorentz violation. In general, the SME may be thought up as a structure which enable scientists to translate the absence of experimentally observed signs of Lorentz violation into constraints on other meaningful physical quantities to be measured [22].
@@
Here we shall be considering a Lorentz non-invariant theory along these lines. Many elementary aspects of the general SME remain somewhat mysterious and are open opportunities for prospective research. Field theories with atypical properties can, even if those field theories are never likely to be physically realized, provide useful theoretical laboratories for understand how quantum theories in general can behave; pure massless quantum electrodynamics in $1+1$ dimensions is unlikely to exist physically, but that did not prevent the Schwinger model [23, 24] from educating us about there may sometimes be found an infinite number of vacua with different topologies. Gauge invariance properties and radiative corrections to the SME effective field theory have, as a result, achieved a plethora of theoretical attentions [25, 26, 27, 28, 29, 30, 31]. The minimal SME, which includes Lorentz-violating field operators of mass dimension two, three, and four will be the domain for our present analysis. Current physical experiments at relatively low energies are useful for testing precisely this low-energy limit of the full SME. Since many low-energy experiments may be made extremely precise, they may be more likely than data collected closer to the Planck scale to provide empirical proof of a violation of the symmetries of the current standard model.
@@
We are setting out now to find extensions of our previous work [32], through examination of major aspects of Lorentz-violating quantum field theories with emphasis on the behaviors of the modified particle momenta and energies [33], and how they affect decay kinematics. Unusual field theory We shall look specifically at three-body decay rates with Lorentz violation, doing computations with a modified outgoing phase space.
@@
Well-developed discussions of certain scattering cross sections, decay rates, and radiation emission rates in the presence of Lorentz symmetry breaking exist in the literature, and kinematic considerations are often of uppermost importance. Many of the modifications to the standard theory that are necessary are described in Ref. [34]; as an example, the authors calculate the cross section for pair annihilation in the limit where particle momentum is greater than mass but still far below the Planck energy. In Ref. [35], the Compton scattering cross section is computed in the presence of a Lorentz- and CPT-breaking background affecting the electron. In that case, since the Lorentz violation breaks the spin degeneracy of the external state, the velocity and phase space factors in the kinematics depend on the particles’ spins; therefore Casimir’s trick for polarization sums does not work for calculating the unpolarized cross section—which means the modified kinematics cause a a fundamental change in how the rate for the process has to be calculated. The cross section needs to be determined using a basis of explicit polarization states for all the incoming and outgoing particles, which can be quite complicated, even with a single isotropic but Lorentz-violating $b_{\mu}$ term added to the action. The scattering cross section differs greatly from the Klein-Nishina formula, and despite the complexity, the calculation can be performed nonperturbatively in $b_{\mu}$, making it possible to observe that the cross section actually diverges at low photon energies—in sharp contrast to the predictions of various low-energy theorems that assume Lorentz invariance. In the electromagnetic sector of the minimal SME, there may exist a Chern-Simons-like Lorentz-violating term. Besides being of intense theoretical interest because of its unusual gauge symmetry properties and their relationship to radiative corrections, this term modifies in the dispersion relations for electromagnetic waves (in a polarization-dependent fashion). This opens up the possibility of a Cerenkov-type effect in vacuum, which is discussed in Ref. [36], which calculates the exact radiation rate using the modified field equations and show that they essentially agree with the a phase space estimate. The phase space estimate is calculated using a standard matrix element, combined with an evaluation of the available outgoing phase space for the single almost-circular polarization mode into which radiation is kinematically allowed. These kinematic considerations can be extended to a broad class of SME coefficients in the charged fermion sector as well [37, 38].
@@
With a similar methodology, we shall here be calculating decay rates in the a simple model of a three-body decay, with a parent particle such as a kaon or $\eta$ meson decaying into three identical pions, in the presence of Lorentz-violating but CPT-preserving $c_{\mu\nu}$ terms that modify the kinematics for the daughters. As well as outlining the total calculation of a total decay rate, we shall look at how the Lorentz violation affects the boundaries of the allowed region of the Dalitz plot for the process.
@@ 3 Decay Phase Space
This will form the integrand of the modified decay integral, with modified quadratic Lorentz-violating dispersion relations visible in the first three Dirac $\delta$-functions. These enforce the energy-momentum relations for each of the three pions in the final state. If the initial unstable particle is at rest, so that $P^{\mu}=(\sqrt{s},\vec{0}\,)$, the three daughter three-momenta line in a plane. We may therefore take the spatial momentum of the first daughter pion to define the $-z$-direction, $p^{\mu}=(E_{p},0,0,-|\vec{p}\,|)$, with the other two daughter trajectories also lying in the $xz$-plane. In the rest frame of the parent particle, $\theta$ is the angle between $\vec{p}$ and $\vec{k}$, and $\alpha$ is the angle between $\vec{l}$ and $\hat{z}$. This configuration is show in figure 1. (Note that because of the anisotropy in $c_{\mu\nu}$, the three-momentum of a particle may point in a slightly different direction from its velocity.)
@@
However, in a case like this one, the isotropic theory is actually too simple. With identical ultrarelativistic daughter particles, the Lorentz-violating modifications in an isotropic theory can be reduced to a simple rescaling of the conventional result. Actually, this is a fairly natural conclusion. The complexity of the $d\Gamma$ in (3) comes from the angular structure, and if there is no anisotropy $c_{0j}=0$ and $c_{jk}$ is proportional to $\delta_{jk}$, which wipes out all the angular structure in (46). If there are no angular dependences, it does not meaningfully matter which directions the daughter pions are moving; all the changes is an overall rescaling of the dependence of the particle energy on the spatial momentum. So we shall instead introduce a different simplifying parameterization of the background tensor $c_{\mu\nu}$.
@@ 3.1 Single-Coefficient Model
×$\times\,\delta\left[(1+c)E_{k}^{2}+c(\sin\theta-\cos\theta)|\vec{k}|E_{k}-(1-c+c\sin\theta\cos\theta)|\vec{k}|^{2}-m^{2}\right]$
×$\times\,\delta\left[(1+c)E_{l}^{2}+c(\sin\alpha+\cos\alpha)|\vec{l}\,|E_{l}-(1-c+c\sin\alpha\cos\alpha)|\vec{l}\,|^{2}-m^{2}\right]$
×$\times\,\delta^{4}(P-p-k-l).$
However, it is possible to simplify this expression even somewhat further while still retaining nontrivial angular anisotropy. The calculation steps and final expression are simplified greatly if we set $c_{13}=c_{31}=0$. A great deal more in-depth analysis of the physical process will be needed if the mixing of the $x$- and $z$-components is to be kept in our calculation. The angular analysis should be made quite complex if this extra dependence is to be considered, so the decay with this term non-zero will be evaluated later. For now we drop this term because without it a simple cancellation arises without angular approximations. This is, we set $c_{\mu\nu}=c\mathbf{M}^{\prime}_{4\times 4}$, where $\mathbf{M}^{\prime}_{4\times 4}$ agrees with $\mathbf{M}_{4\times 4}$, except that $\left(\mathbf{M}^{\prime}_{4\times 4}\right)_{13}=\left(\mathbf{M}^{\prime}_{4\times 4}\right)_{31}=0$. With this modification, the differential decay rate for the channel is
@@
×$\times\,\delta\left[(1+c)E_{k}^{2}+c(\sin\theta-\cos\theta)|\vec{k}|E_{k}-(1-c)|\vec{k}|^{2}-m^{2}\right]$
×$\times\,\delta\left[(1+c)E_{l}^{2}+c(\sin\alpha+\cos\alpha)|\vec{l}\,|E_{l}-(1-c)|\vec{l}\,|^{2}-m^{2}\right]\,\delta^{4}(P-p-k-l).$
@@
×$\times\,\delta\left[(1+c)E_{k}^{2}+c(\sin\theta-\cos\theta)|\vec{k}|E_{k}-(1-c)|\vec{k}|^{2}-m^{2}\right]$
×$\times\,\delta\left[\sqrt{s}-E_{p}-E_{k}-E_{l}(\vec{l}=-\vec{p}-\vec{k}\,)\right].$
Here, ±$\sum_{\pm}\delta(E_{l}-\omega_{3}^{\pm})$ appears because of the mathematical possibility of $E_{l}$, which is determined by a quadratic energy-momentum relation, could potentially takes roots $\omega_{l}^{\pm}(\vec{l}\,)$ of either positive or negative sign. Physically, however, obviously only the positive-energy root
@@
contributes to the the decay rate.
@@
$\delta(\Delta)$ $=$ $\delta\left(\sqrt{s}-\frac{1}{2(1+c)}\left\{\sqrt{|\vec{p}\,|^{2}(4-3c^{2})+4(1+c)m^{2}}\right.\right.$ (15)
$\left.\left.+\sqrt{|\vec{k}|^{2}(4-3c^{2}-2c^{2}\cos\theta\sin\theta)+4(1+c)m^{2}}\right.\right.$
@@
After dividing the integrand in $d\Gamma$ by $|\partial\Delta/\partial\theta|$, the resulting integral is manageable at first order in $c$. The terms which depend on $\theta$ cancel after the term is divided into the integral and then the whole integrand approximated to first order in $c$. This avoids the complicated integration that would be necessary if the angular roots needed be plugged back into $\theta$-dependent terms remaining in the integrand.
At this point, it is not possible to go further using only facts about the outgoing particle kinematics. For a generic process, the matrix element $i\mathcal{M}$ will also be a function of $\vec{p}$ and $\vec{k}$ (as well as the constrained momentum $\vec{l}$. The dependence can be both on the magnitudes of the three outgoing pion momenta and the angles between them. For the purpose of continuing our kinematical analysis, we must therefore make choice of what matrix element to use. To simplify things to the greatest extent possible—so as to keep the focus on what effects may be attributed purely to daughter particle kinematics—we shall opt for the simplest possible choice: a constant $|\mathcal{M}|^{2}$, such as might arise from a four-meson contant interaction of the form $\mathcal{L}_{I}\propto\lambda(\eta\pi^{3})$. Henceforth, the $|\mathcal{M}|^{2}$ appearing in our expressions will be treated as a momentum-independent constant. This constant may generically include additional factors of $c$; in particular, the rescaling of the $\pi$ field to remove $c_{00}\partial^{0}\partial^{0}$ terms mentioned above can introduce $\mathcal{O}(c)$ momentum-independent corrections to the matrix element.
@@
With the angular dependence of the matrix element squared known (and in our toy model case, trivial), it is possible to perform the angular integrations explicitly, at least to leading order in the Lorentz violation. The integration over the angular variables that the integrand does not depend upon (meaning over the direction of $\vec{p}$, and the azimuthal angle of $\vec{k}$ produces a factor a factor of $2(2\pi)^{2}$. The last angular integration over $\theta$ gives new factors that yield
@@
Although we have not yet integrated over the magnitudes of the spatial momenta [or, equivalently, over $p^{(0)}$ and $k^{(0)}$], but only the associated angles, any momentum-dependent matrix element $i\mathcal{M}$ would need to by symmetric in $\vec{p}$, $\vec{k}$, and $\vec{l}=-\vec{p}-\vec{k}$; and the symmetric dependence on $\vec{l}$ means that if $i\mathcal{M}$ depends on the magnitudes of the three-momenta, it will also unavoidably depend on $\theta$. So only with a completely constant matrix element may we reach this and later expressions.
@@
[…]
@@
[…]

### v2

# arXiv:2506.19564

@@ 1 Introduction
In the standard model, physics must be identical in each inertial frame: invariant under Lorentz transformations. However, we relax this notion and implement particle theory models in inertial frames where the physics is modified. These Lorentz noninvariant theories contain inertial frames that are not all equivalent. This makes necessary changes to specific calculation procedures used in the standard model; we can obtain completely or slightly modified observable values or predict new physical phenomena that can also be measured. For example, changes in the standard energy-momentum relations for various types of quanta can affect the rates and thresholds for particle processes.
@@
Theories with Lorentz symmetry breaking are also of interest, since quantum theories of gravitation may not have the same symmetries as the low-energy effective theories that we see in everyday operation. Theories of Planck-scale physics like strings, loop quantum gravity, noncommutative spacetime structures, spacetime foam, nontrivial spacetime topologies, and others already predict or strongly suggest Lorentz violation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. The same tools used to study the standard model may also be used to analyze more general theories in which Lorentz symmetry is weakly broken. The general local field theory that describes Lorentz-violating modifications to processes involving known standard model fields has been laid out in detail [15, 16]. This theory, which is called the standard model extension (SME), is also the general theory for describing local CPT violation involving those known fields (although, while the SME may be extended to include gravity [17], gravitational CPT violation is a more complicated matter [18, 19, 20]); the reason is that a local, stable, unitary, CPT-violating quantum field theory is also necessarily Lorentz violating [21]. The SME was constructed precisely to provide a systematic approach, via an effective field theory framework for understanding searches for violations of the fundamental Lorentz and CPT symmetries.
@@
As the SME is an effective field theory, it contains towers of operators of increasing mass dimension, with the contributions from higher-dimensional operators assumed to be further suppressed under most circumstances. The minimal SME is a restricted subtheory of the SME—that which is expected to be renormalizable, because it satisfies the usual conditions that ensure renormalizability for the standard model, by containing only the finite number of local, Hermitian, gauge-invariant operators which can be built out of the known standard model fermion and boson field that are of dimension four or less. The terms in the minimal SME Lagrange density are, in structure, very similar to those appearing the usual standard model; the key difference is that the additional Lorentz-violating operators each have one or more uncontracted Lorentz indices. The coefficients multiplying those operators then become observable as components of preferred background vectors or tensors. In many situations, the minimal SME is the most natural test theory for analyzing the results of experimental searches for CPT or Lorentz violation. In general, the SME may be thought up as a structure which enable scientists to translate the absence of experimentally observed signs of Lorentz violation into constraints on other meaningful physical quantities to be measured [22].
@@
Here we shall be considering a Lorentz noninvariant theory along these lines. Many elementary aspects of the general SME remain somewhat mysterious and are open opportunities for prospective research. Field theories with atypical properties can, even if those field theories are never likely to be physically realized, provide useful theoretical laboratories for understanding how quantum theories in general can behave; pure massless quantum electrodynamics in $1+1$ dimensions is unlikely to exist physically, but that did not prevent the Schwinger model [23, 24] from educating us about the fact that there may sometimes be an infinite number of vacua with different topologies. Gauge invariance properties and radiative corrections to the SME effective field theory have, as a result, achieved a plethora of theoretical attentions [25, 26, 27, 28, 29, 30, 31]. The minimal SME, which includes Lorentz-violating field operators of mass dimension two, three, and four will be the domain for our present analysis. Current physical experiments at relatively low energies are useful for testing precisely this low-energy limit of the full SME. Since many low-energy experiments may be made extremely precise, they may be more likely than data collected closer to the Planck scale to provide empirical proof of a violation of the symmetries of the current standard model.
@@
We are setting out now to find extensions of our previous work [32], through examination of major aspects of Lorentz-violating quantum field theories with emphasis on the behaviors of the modified particle momenta and energies [33], and how they affect decay kinematics in an unusual field theory. We shall look specifically at three-body decay rates with Lorentz violation, doing computations with a modified outgoing phase space.
@@
Well-developed discussions of certain scattering cross sections, decay rates, and radiation emission rates in the presence of Lorentz symmetry breaking exist in the literature, and kinematic considerations are often of uppermost importance. Many of the modifications to the standard theory that are necessary are described in Ref. [34]; as an example, the authors calculate the cross section for pair annihilation in the limit where particle momentum is greater than mass but still far below the Planck energy. In Ref. [35], the Compton scattering cross section is computed in the presence of a Lorentz- and CPT-breaking background affecting the electron. In that case, since the Lorentz violation breaks the spin degeneracy of the external state, the velocity and phase space factors in the kinematics depend on the particles’ spins; therefore Casimir’s trick for polarization sums does not work for calculating the unpolarized cross section—which means the modified kinematics cause a fundamental change in how the rate for the process has to be calculated. The cross section needs to be determined using a basis of explicit polarization states for all the incoming and outgoing particles, which can be quite complicated, even with a single isotropic but Lorentz-violating $b_{\mu}$ term added to the action. The scattering cross section differs greatly from the Klein-Nishina formula, and despite the complexity, the calculation can be performed nonperturbatively in $b_{\mu}$, making it possible to observe that the cross section actually diverges at low photon energies—in sharp contrast to the predictions of various low-energy theorems that assume Lorentz invariance. In the electromagnetic sector of the minimal SME, there may exist a Chern-Simons-like Lorentz-violating term. Besides being of intense theoretical interest because of its unusual gauge symmetry properties and their relationship to radiative corrections, this term modifies in the dispersion relations for electromagnetic waves (in a polarization-dependent fashion). This opens up the possibility of a Cerenkov-type effect in vacuum, which is discussed in Ref. [36], which calculates the exact radiation rate using the modified field equations and shows that they essentially agree with the phase space estimate. The phase space estimate is calculated using a standard matrix element, combined with an evaluation of the available outgoing phase space for the single almost-circular polarization mode into which radiation is kinematically allowed. These kinematic considerations can be extended to a broad class of SME coefficients in the charged fermion sector as well [37, 38].
@@
With a similar methodology, we shall here be calculating decay rates in a simple model of a three-body decay, with a parent particle such as a kaon or $\eta$ meson decaying into three identical pions, in the presence of Lorentz-violating but CPT-preserving $c_{\mu\nu}$ terms that modify the kinematics for the daughters. As well as outlining the calculation of a total decay rate, we shall look at how the Lorentz violation affects the boundaries of the allowed region of the Dalitz plot for the process.
@@ 2 Lorentz-Violating Field Theory
In this paper, she shall focus on two specific textures for the $c_{\mu\nu}$ background tensor. A more complete future analysis, covering all forms of anisotropy that could arise from a scalar-sector $c_{\mu\nu}$, would consider sets of coefficients having different symmetry types separately, to see numerically how those of each individual type would shift the shape of the Dalitz plot for a specific decay at a known energy. When possible, bounds on SME coefficients are typically reported in a standardized celestial equatorial coordinate system. Time-dependent coordinate transformations may then be used to relate time-stamped laboratory data to coefficients in the standardized coordinates. The coefficients in celestial equatorial coordinates will determine the amplitudes for time-dependent oscillations in the shapes of the plots, at the Earth’s sidereal rotation frequency $\omega_{\oplus}$ and also at $2\omega_{\oplus}$ [39].
@@ 3 Decay Phase Space
This will form the integrand of the modified decay integral, with modified quadratic Lorentz-violating dispersion relations visible in the first three Dirac $\delta$-functions. These enforce the energy-momentum relations for each of the three pions in the final state. If the initial unstable particle is at rest, so that $P^{\mu}=(\sqrt{s},\vec{0}\,)$, the three daughter three-momenta lie in a plane. We may therefore take the spatial momentum of the first daughter pion to define the z-z-direction, $p^{\mu}=(E_{p},0,0,-|\vec{p}\,|)$, with the other two daughter trajectories also lying in the $xz$-plane. In the rest frame of the parent particle, $\theta$ is the angle between $\vec{p}$ and $\vec{k}$, and $\alpha$ is the angle between $\vec{l}$ and $\hat{z}$. This configuration is shown in figure 1. (Note that because of the anisotropy in $c_{\mu\nu}$, the three-momentum of a particle may point in a slightly different direction from its velocity.)
@@
However, in a case like this one, the isotropic theory is actually too simple. With identical ultrarelativistic daughter particles, the Lorentz-violating modifications in an isotropic theory can be reduced to a simple rescaling of the conventional result. Actually, this is a fairly natural conclusion. The complexity of the $d\Gamma$ in (3) comes from the angular structure, and if there is no anisotropy $c_{0j}=0$ and $c_{jk}$ is proportional to $\delta_{jk}$, which wipes out all the angular structure in (46). If there are no angular dependences, it does not meaningfully matter which directions the daughter pions are moving; the only change is an overall rescaling of the dependence of the particle energy on the spatial momentum. So we shall instead introduce a different simplifying parameterization of the background tensor $c_{\mu\nu}$.
@@ 3.1 Single-Coefficient Model
$\times\,\delta\left[(1+c)E_{k}^{2}+c(\sin\theta-\cos\theta)|\vec{k}|E_{k}-(1-c+c\sin\theta\cos\theta)|\vec{k}|^{2}-m^{2}\right]$
$\times\,\delta\left[(1+c)E_{l}^{2}+c(\sin\alpha+\cos\alpha)|\vec{l}\,|E_{l}-(1-c+c\sin\alpha\cos\alpha)|\vec{l}\,|^{2}-m^{2}\right]$
$\times\,\delta^{4}(P-p-k-l).$
However, it is possible to simplify this expression even somewhat further while still retaining nontrivial angular anisotropy. The calculation steps and final expression are simplified greatly if we set $c_{13}=c_{31}=0$. A great deal more in-depth analysis of the physical process will be needed if the mixing of the $x$- and $z$-components is to be kept in our calculation. The angular analysis should be made quite complex if this extra dependence is to be considered, so the decay with this term nonzero will be evaluated later. For now we drop this term because without it a simple cancellation arises without angular approximations. This is, we set $c_{\mu\nu}=c\mathbf{M}^{\prime}_{4\times 4}$, where $\mathbf{M}^{\prime}_{4\times 4}$ agrees with $\mathbf{M}_{4\times 4}$, except that $\left(\mathbf{M}^{\prime}_{4\times 4}\right)_{13}=\left(\mathbf{M}^{\prime}_{4\times 4}\right)_{31}=0$. It is common practice, when determining the maximum sensitivity of searches for Lorentz violation, to consider only restricted subsets of SME coefficients, rather than all of them at once. Operators with different parities under discrete symmetries will not mix under quantum corrections or through other modes of mediation. Since $c_{13}$ has different behavior under time reversal, charge conjugation, and spatial reflections along the $x$- and $z$-axes from any other minimal SME operator, it may be safely omitted. If $c_{13}=0$ in the fundamental action, it will be protected from being generated by any indirect effects.
With this modification, the differential decay rate for the channel is
@@
$\times\,\delta\left[(1+c)E_{k}^{2}+c(\sin\theta-\cos\theta)|\vec{k}|E_{k}-(1-c)|\vec{k}|^{2}-m^{2}\right]$
$\times\,\delta\left[(1+c)E_{l}^{2}+c(\sin\alpha+\cos\alpha)|\vec{l}\,|E_{l}-(1-c)|\vec{l}\,|^{2}-m^{2}\right]\,\delta^{4}(P-p-k-l).$
@@
$\times\,\delta\left[(1+c)E_{k}^{2}+c(\sin\theta-\cos\theta)|\vec{k}|E_{k}-(1-c)|\vec{k}|^{2}-m^{2}\right]$
$\times\,\delta\left[\sqrt{s}-E_{p}-E_{k}-E_{l}(\vec{l}=-\vec{p}-\vec{k}\,)\right].$
Here, $\sum_{\pm}\delta(E_{l}-\omega_{3}^{\pm})$ appears because of the mathematical possibility of $E_{l}$, which is determined by a quadratic energy-momentum relation, could potentially takes roots $\omega_{l}^{\pm}(\vec{l}\,)$ of either positive or negative sign. Physically, however, obviously only the positive-energy root
@@
contributes to the decay rate.
@@
[…]
@@
After dividing the integrand in $d\Gamma$ by $|\partial\Delta/\partial\theta|$, the resulting integral is manageable at first order in $c$. The terms which depend on $\theta$ cancel after the term is divided into the integral and then the whole integrand approximated to first order in $c$. This avoids the complicated integration that would be necessary if the angular roots needed to be plugged back into $\theta$-dependent terms remaining in the integrand.
At this point, it is not possible to go further using only facts about the outgoing particle kinematics. For a generic process, the matrix element $i\mathcal{M}$ will also be a function of $\vec{p}$ and $\vec{k}$ (as well as the constrained momentum $\vec{l}$). The dependence can be both on the magnitudes of the three outgoing pion momenta and the angles between them. For the purpose of continuing our kinematical analysis, we must therefore make choice of what matrix element to use. To simplify things to the greatest extent possible—so as to keep the focus on what effects may be attributed purely to daughter particle kinematics—we shall opt for the simplest possible choice: a constant $|\mathcal{M}|^{2}$, such as might arise from a four-meson contact interaction of the form $\mathcal{L}_{I}\propto\lambda(\eta\pi^{3})$. Henceforth, the $|\mathcal{M}|^{2}$ appearing in our expressions will be treated as a momentum-independent constant. This constant may generically include additional factors of $c$; in particular, the rescaling of the $\pi$ field to remove $c_{00}\partial^{0}\partial^{0}$ terms mentioned above can introduce $\mathcal{O}(c)$ momentum-independent corrections to the matrix element.
@@
With the angular dependence of the matrix element squared known (and in our toy model case, trivial), it is possible to perform the angular integrations explicitly, at least to leading order in the Lorentz violation. The integration over the angular variables that the integrand does not depend upon (meaning over the direction of $\vec{p}$, and the azimuthal angle of $\vec{k}$) produces a factor a factor of $2(2\pi)^{2}$. The last angular integration over $\theta$ gives new factors that yield
@@
Although we have not yet integrated over the magnitudes of the spatial momenta [or, equivalently, over $p^{(0)}$ and $k^{(0)}$], but only the associated angles, any momentum-dependent matrix element $i\mathcal{M}$ would need to be symmetric in $\vec{p}$, $\vec{k}$, and $\vec{l}=-\vec{p}-\vec{k}$; and the symmetric dependence on $\vec{l}$ means that if $i\mathcal{M}$ depends on the magnitudes of the three-momenta, it will also unavoidably depend on $\theta$. So only with a completely constant matrix element may we reach this and later expressions.
@@
$d\Gamma=\frac{1}{64\pi^{2}}\frac{|\mathcal{M}|^{2}}{\sqrt{s}}\,dp^{(0)}\int\limits_{\frac{\sqrt{s}(1+c)}{2}-p^{(0)}}^{\frac{\sqrt{s}(1+c)}{2}}dk^{(0)}\left\{1+c-\frac{c}{2}\left[\frac{m^{2}}{(p^{(0)})^{2}}+\frac{m^{2}}{(k^{(0)})^{2}}\right]\right\}.$ (21)
@@
$\Gamma=\frac{1}{64\pi^{2}}\frac{|\mathcal{M}|^{2}}{\sqrt{s}}\int\limits_{m}^{\frac{\sqrt{s}(1+c)}{2}}dp^{(0)}\left[p^{(0)}+c\left(p^{(0)}-\frac{m^{2}}{2p^{(0)}}-\frac{2p^{(0)}m^{2}}{s-2\sqrt{s}p^{(0)}}\right)\right].$ (22)